If the modulus of elasticity of an elliptical hair is 3.75 GPa. The dimensions of the hair are A = 40μ, B=30 μ. The stress at which a hair breaks is called the breaking stress. Suppose this hair broke under a load of 88 grams. What was the breaking stress?
Answer #1

Applied load=88 gms=0.088*9.8=0.86 N

Area of sueface on which the load is applied=(40*10^-6)*(30*10^-6)=1.2*10^-9 m²

Stress= force /area=0.86/(1.2*10^-9)=0.717*10^9 Pa=0.717 GPa

This is the required answer